OpenStudy (haltiamali):

(2a^3b^2) (b^3)^2 Simplify. (The ^'s mean to the power of..)

OpenStudy (ajprincess):

\(\large (x^y)^z)=x^{y*z}\) using this can u simplify \(\large (b^3)^2\)?

OpenStudy (haltiamali):

So it would be \[b^6..?\]

OpenStudy (ajprincess):

ya very good:)

OpenStudy (ajprincess):

\(\large x^y*x^z=x^{y+z}\) Usin this can u simplify \(\large b^2*b^6\)?

OpenStudy (haltiamali):

Is the answer \[2a^5b^12\]

OpenStudy (haltiamali):

that last number is supposed to be a 12

OpenStudy (ajprincess):

nope. \(\large x^y*x^z=x^{y+z}\) using this hint can u simplify \(\large b^2*b^6\)?

OpenStudy (haltiamali):

\[b^8?\]

OpenStudy (ajprincess):

ya very good:)

OpenStudy (haltiamali):

\[2a^3b^8\] is the answer?

OpenStudy (ajprincess):

yup:)

OpenStudy (haltiamali):

thx. :) can you help me with the other 3?

OpenStudy (ajprincess):

ya sure:)

OpenStudy (haltiamali):

\[(-4xy)^3 (-2x^2)^3\]

OpenStudy (ajprincess):

\(\Large (-2x^2)^2=(-2)^2*(x^2)^2\) What is \(\Large (-2)^2\)?

OpenStudy (haltiamali):

4

OpenStudy (ajprincess):

ya good:) What do u get when u simplify \(\Large (x^2)^2\)?

OpenStudy (haltiamali):

4x?

OpenStudy (ajprincess):

nope \(\Large (a^b)^c=a^{b*c}\) using this simplify \(\Large (x^2)^2\)

OpenStudy (haltiamali):

ohh. \[x^4\] right?

OpenStudy (ajprincess):

ya:) \(\Large(-4xy)^3=(-4)^3*(x^1)^3*(y^1)^3\) Nw can u simplify (-4)^3*(x^1)^3*(y^1)^3\)

OpenStudy (haltiamali):

-64 x^4 y^4?

OpenStudy (haltiamali):

oh wait

OpenStudy (haltiamali):

hahaha I get it :D

OpenStudy (haltiamali):

\[(-14xy^4) (-1/2xy^3)\]

OpenStudy (haltiamali):

that's negative half

OpenStudy (haltiamali):

Would it be \[-14 1/2 x+y^7?\]

OpenStudy (ajprincess):

wait a sec. oops sorry it is \(\Large (-2x^2)^3\) right? So \(\Large (-2x^2)^3=(-2)^3*(x^2)^3\) We r nt yet done with our previous problem. so it will be better if we finish that and come to this one. What is \((-2)^3)? What do u get when u simplify \((x^2)^3\)?

OpenStudy (haltiamali):

\[(-2x^6) \] right?

OpenStudy (ajprincess):

\(\Large (x^2)^3=x^6\) that's right:) but what is \(\Large (-2)^3=?

OpenStudy (haltiamali):

oh -8

OpenStudy (ajprincess):

ya:) so \(\Large (-2x^2)^3=-8x^6\) Nw \(\Large(-4xy)^3=(-4)^3*(x^1)^3*(y^1)^3\) What is \(\Large (-4)^3=\)? What do u get when u simplify \(\Large (x^1)^3\) and \(\Large (y^1)^3\)?

OpenStudy (haltiamali):

-64 x^4 and y^4

OpenStudy (ajprincess):

\(\Large(a^b)^c=a^{b*c}\) usin this general rule only u have to simplify \((x^1)^3\) and \((y^1)^3\)

OpenStudy (haltiamali):

oh. x^3 and y^3

OpenStudy (ajprincess):

ya good:)

OpenStudy (ajprincess):

\(\Large(-4xy)^3=-64x^3y^3\) \(\Large(-2x2)^3=-8x^6\) \(\Large(-64x^3y^3)*(-8x^6)=(-64*8)*(x^3*x^6)*y^3\) To simplify \(\Large(x^3*x^6)\) u have to use the general rule \(\Large(a^b)*(a^c)=a^{b+c}\)

OpenStudy (ajprincess):

Can u figure out the final answer @haltiamali

OpenStudy (haltiamali):

yes :) tysm

OpenStudy (ajprincess):

k nw moving on to ur next problem \[(-14xy^4)(\frac{ -1 }{ 2 }xy^3)=(-14*-\frac{ 1 }{ 2 })(x^1*x^1)*(y^4*y^3)\] See if u can simplify nw. If u find any difficulties while solving pls do ask:)

OpenStudy (haltiamali):

Thanks ! :)

OpenStudy (ajprincess):

Welcome:)